a cylinder with a diameter of 22 inches and a volume of 15,19706 cu. in. find the depth of the cylinder

Answer by lwsshak3(11628)

a cylinder with a diameter of 22 inches and a volume of 15,19706 cu. in. find the depth of the cylinder
**
volume=area of end*depth=1519706
d=diameter
area of end=πd^2/4
πd^2/4*depth=1519706
depth=1519706/π22^2/4
≈3997.83 in